A golf ball with a mass of 0, 1 kg is driven from the tee. The golf ball experiences a force of 1000 N while in contact with the golf club and moves away from the golf club at 30 m·s-1. Calculate the time during which the golf club was in contact with the ball. 


1
Expert's answer
2022-05-19T16:40:20-0400

The second Newton's law says


"F=\frac{\Delta p}{\Delta t}"

Thus, the time during which the golf club was in contact with the ball

"\Delta t=\frac{m\Delta v}{F}=\frac{0.1*30}{1000}=0.003\:\rm s"